Find the eccentricity and coordinates of the foci of ellipse. X2+3y2=a2
 
 
 

Grade:11

2 Answers

Arun
5 years ago
 
Dear student
 
we can write it as-
x^2 /a^2.  + y^2 /(a^2/3)= 1
 
hence eccentricity
 
b^2 =a^2(1-e^2)
 
1/3 = 1 – e^2
 
e^2 = 2/3
 
e = sqrt (2/3)
Deepak Kumar Shringi
5 years ago
semi major axis = a
semi minor axis= (a/3)1/2
e=(2/3)1/2
focus = a(2/3)1/2